If the side of a square is increased by 4 cm, the perimeter becomes 136 cm. Find the length of the side of the original square.
Answer by mananth(16949)

let side of square be x
it is increased by 4
new length = (x+4)
Perimeter of square = 4*L
=4*(x+4)
=4x+16
..
4x+16= 136
add -16 to both sides
4x+16-16=136-16
4x=120
divide by 4
4x/4 = 120/4
x=30 cm the side of original square
..